It costs $24 per hour plus a flat fee of $23 for a plumber to make a house call. After writing an equation for this situation, what is the total cost to have a plumber come to a house for 10 hours?
Answer by lwsshak3(11628)

It costs $24 per hour plus a flat fee of $23 for a plumber to make a house call. After writing an equation for this situation, what is the total cost to have a plumber come to a house for 10 hours?
..
You can think of this relationship as a straight line applying the standard form, y=mx+b, with m=slope and b=y-intercept.
Picture a straight line graph with cost on the y-axis and hours on the x-axis. At zero hours, just when the plumber enters the house, the cost is already $23, from hereon, every hour costs an additional $24. This is the slope in the equation which is as follows:
y=24x+23
Total cost for 10 hours=24*10+23=$263